OpenStudy (anonymous):

quick question: MEDAL + FAN - - how is 'twice the number of students in juans class divided by ten' written as a equation? [ use j as juan ]

OpenStudy (abbles):

2J/10 Where 2 and J are being multiplied and then divided by 10. Oh and I used 'J' as the number of students in Juan's class.

OpenStudy (anonymous):

@Abbles yeah thats what I meant and wouldn't it be 2j/5 = 10? or nah

OpenStudy (anonymous):

@Abbles or 2j/5 + 10?

OpenStudy (abbles):

No because you would need to divide the number by ten, not add. Reread the question and maybe you'll see... it says 'divided by ten' at the end. So if that is written correctly it should be: 2J/10 You could also write it: 2J/10 = x x being the number you are trying to figure out :)

OpenStudy (anonymous):

@Abbles in my answer choices it was 2j/5 = 10 and it was right (:

OpenStudy (abbles):

Really? Are you sure you wrote the question correctly and all of it? That is so weird. 'twice the number of students in juans class divided by ten' should be 2J/10 because that's all it's saying... unless you left something out? anyway I'm glad you got the correct answer!

OpenStudy (anonymous):

@Abbles yeah I swear I wish I could show you

OpenStudy (anonymous):

@Abbles either a typo by who wrote it or my mistake haha

OpenStudy (abbles):

That is bizarre but hey, at least you got the right answer! good job :)

OpenStudy (anonymous):

@Abbles haha yeah thanks so much :D

OpenStudy (abbles):

Yeah, anytime :) Let me know if you have any other questions.